# Flow Matching Posterior Sampling: A Training-free Conditional Generation for Flow Matching

### v2

@@ 2.3 Training-free Conditional Generation based on FMs
Several approaches have been proposed to address this issue. For instance, D-Flow [10] introduced a lightweight framework that fine-tunes the initial point of the generation process using a given condition, effectively guiding the trajectory toward $p(x|c)$. In contrast, $\pi$Flow [16] first exploits the theoretical relationship between FMs and SDMs to re-establish a formulation of $p(x|c)$ based on $p(x)$, and then proposes a closed-form solution under the assumption of linear inversion. Feng et al. [17] developed a novel theoretical framework that directly links $p(x|c)$ with the distribution $p(x)$. This approach leads to a closed-form expression for $p(x|c)$, which, however, is analytically intractable. To overcome this limitation, the authors further apply the mentor-color theory to approximate the intractable form, again relying on the linear inversion assumption.
@@ 3.2 Flow Matching Models
where $h$ is the time interval between the $t$-th and $(t-1)$-th time steps.
@@ 4.1 The Condition-related Correction Term
It could be found that Eq 21 and Eq. 12 could be matched by $\Delta(x_{t},c)=-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})$. Then, $\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})$ exactly represents the direction from $x_{t}$ to $c$. Thus, we could use $-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})$ to formulate the correction term by:
@@ 4.2 FMPS
$g^{1}=\frac{||v(x_{t},t)||_{2}}{||\nabla_{x_{t}}D(\hat{x}_{0|t},c)||_{2}}\nabla_{x_{t}}D(\hat{x}_{0|t},c).$ (28)
Eq. 28 limits the strength of $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ under the bound of $||v(x_{t},t)||_{2}$, where $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ will not too strong and too weaken.
@@
Figure 2: Quantitative comparison of linear/nonlinear tasks on CelebA-HQ and AFHQ [16]. For linear tasks, we mainly report the results between FMPS and $\Pi$FLow, which is the SOTA method for FMs. For nonlinear tasks, we mainly report the results between FMPS and FreeDom. For FreeDom, we use the optimal setting from UFG [34].
@@ 5.1 Implementation Details
5.2 Distance Metric
All the distance metrics $D(\ast,\ast)$ we use are the same as those in previous works.
5.2.1 Linear Inverse Problems
Inpainting. Given a known center box inpainting mask $m_{i}$ and the Gaussian noise $\epsilon$, the measurement $y$ could be represented as:
$y=m_{i}\odot x+0.05\epsilon,$ (31)
where $x$ is the unknown image. Then, the $d(\ast,\ast)$ is defined as:
$D(\hat{x}_{0|t},y)=||(m_{i}\odot\hat{x}_{0|t}+0.05\epsilon)-y||_{2}.$ (32)
Super-resolution. Given a known interpolation-based super-resolution operator $f_{sr}(\ast)$ (e.g., 4x upsampling), and additive Gaussian noise $\epsilon$, the measurement $y$ can be modeled as:
$y=f_{sr}(x)+0.05\epsilon.$ (33)
Then, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||(f_{sr}(\hat{x}_{0|t})+0.05\epsilon)-y||_{2}.$ (34)
Gaussian Deblur. Given a known Gaussian blur operator $f_{gb}(\ast)$ with 61 kernel size and 3 intensify and the Gaussian noise $\epsilon$, the measurement $y$ could be represented as:
$y=f_{gb}(x)+0.05\epsilon.$ (35)
Then, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||(f_{gb}(\hat{x}_{0|t})+0.05\epsilon)-y||_{2}.$ (36)
5.2.2 Non-linear Inverse Problems
Segmentation. Given a pre-trained face segmentation model $f_{seg}$, the measurement $y$, where the output of $f_{seg}$ is the face segmentation mask map and $y$ is random known segmentation map, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||f_{seg}(\hat{x}_{0|t})-y||_{2}.$ (37)
Sketch. Given a pre-trained image to sketch model $f_{ske}$, the measurement $y$, where the output of $f_{ske}$ is the sketch related to the input image and $y$ is random known sketch, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||f_{ske}(\hat{x}_{0|t})-y||_{2}.$ (38)
Face ID Given a pre-trained face classifier $f_{face}$, the measurement $y$, where the output of $f_{face}$ is the logit value related to the input images and $y$ is random known image, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||f_{face}(\hat{x}_{0|t})-f_{face}(y)||_{2}.$ (39)
5.2.3 Text-Style generation
Given the image encoder from a pre-trained CLIP [43] noted as $E_{i}$ and the style image $y$, the $D(\ast,\ast)$ is:
$D(\hat{x}_{0|t},y)=||\text{Gram}(E_{i}(\hat{x}_{0|t}))-\text{Gram}(E_{i}(y))||_{2},$ (40)
where $\text{Gram}(\ast)$ is the gram matrix operator [12].
@@ 5.3 Qualitative Results
Text-style generation. To prove that the FMPS could work on the real generation tasks, we first report the comprehensive qualitative evaluation in Table 4 using 1000 text-image pairs under nine style types. It can be observed that FMPS-free more closely adheres to the text prompt guidelines, achieving the highest CLIP score while maintaining a low style distance. Then, to show the ability of FMPS-free, we report the qualitative results using the complex text prompts shown in Fig. LABEL:fig:style. It can be noticed that FMPS-free achieves a better generation since it maintains a similar style, following the style of images while satisfying the text prompts.
@@ 5.4 Quantitative Results
We report the comprehensive quantitative results for both linear/nonlinear (shown in Fig. 2) and style text-to-image (shown in Fig. 3). It could be found that FMPS outperforms the previous methods in all tasks.

### v3

@@ 2.3 Training-free Conditional Generation based on FMs
[…]
@@ 3.2 Flow Matching Models
where $h$ is the time interval between the $t$-th and (t1)(t-1)-th time steps.
@@ 4.1 The Condition-related Correction Term
It could be found that Eq 21 and Eq. 12 could be matched by $\Delta(x_{t},c)=-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})$. Then, $\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})$ exactly represents the direction from $x_{t}$ to $c$. Thus, we could use $\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})$ to formulate the correction term by:
@@ 4.2 FMPS
[…]
Eq. 28 limits the strength of $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ under the bound of $v(x_{t},t)||_{2}$, where $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ will not too strong and too weaken.
@@ 5.1 Implementation Details
Distance metric. The distance metric used in this paper is the same as the previous works [12]. The details are in the supplementary.
@@ 5.2 Qualitative Results
Text-style generation. To prove that the FMPS could work on the real generation tasks, we first report the comprehensive qualitative evaluation in Table 4 using 1000 text-image pairs under nine style types. It can be observed that FMPS-free more closely adheres to the text prompt guidelines, achieving the highest CLIP score while maintaining a low style distance. Then, to show the ability of FMPS-free, we report the qualitative results using the complex text prompts shown in Fig. 2. It can be noticed that FMPS-free achieves a better generation since it maintains a similar style, following the style of images while satisfying the text prompts.
@@ 5.3 Quantitative Results
We report the comprehensive quantitative results for both linear/nonlinear (shown in the supplementary) and style text-to-image (shown in Fig. 2). It could be found that FMPS outperforms the previous methods in all tasks.